# Prediction Markets with Intermittent Contributions

### v1

@@ I Introduction
A solution to increase collaboration is to give incentives in exchange for these data. This can be done in different ways, depending on whether the companies are willing to share data. If this is the case, different methods have been proposed for renewable power forecasting in the literature exploiting spatio-temporal relations. For solar power forecasting, a model incorporating exogenous data from different sources was developed [1], while a similar approach has been applied to both wind and solar power forecasting [2]. In wind power forecasting, sparse spatio-temporal models were introduced in [3] and later extended in an online learning context [4]. Furthermore, to avoid the loss of privacy companies could face, privacy-preserving solutions using distributed learning methods have been proposed [5], with further advancements addressing online settings [6, 7]. However, all these solutions assume that the agents are willing to collaborate to improve forecasts and that they act rationally and truthfully. This is not always the case in practice. An alternative approach is to view the problem in a more general framework, through prediction markets, in which companies choose to share their individual forecasts and get rewarded for their contribution to the resulting aggregate forecast.
@@
Nevertheless, existing solutions fail to consider some essential aspects of real-world applications such as: (i) real-time implementation, (ii) historical contributions of the participants, and (iii) the ability to accommodate intermittent participation, meaning that participants may join or exit the market at any time. The first two challenges can be seen as an online learning problem, whilst the third one as a missing data problem. For the latter, the literature has proposed robust variants of linear regression [11, 12] and online imputation strategies [13]. In addiction, a pay-off allocation mechanism must be designed to determine how rewards are distributed among sellers, reflecting the informational value they contribute to the final forecast. Also in this case, the allocation must reflect historical contributions rather than only instantaneous performance. In the literature, the most widely adopted approach is the Shapley-value-based allocation [14] (see, for example, [7]) and online versions of Shapley value calculation have also been employed for dynamic pay-off distribution [15]. However, these approaches primarily focus on in-sample pay-offs and fail to account for out-of-sample performance, overlooking the actual contribution of sellers to genuine forecast improvements. Finally, any allocation rule must also satisfy several desirable economic properties to ensure fairness and stability.
@@ IV-B Performance with Varying Missingness
RQR
$4.8\pm 3.9$ $1.7\pm 0.14$ $5\pm 6.5$ $1.7\pm 0.2$ $0.3\pm 4$ $1.7\pm 0.16$
Last
Impute
$5\pm 4.6$ $2.5\pm 0.2$ $35\pm 7$ $2.5\pm 0.3$ $18\pm 6$ $2.5\pm 0.2$
Mean
Impute
$7\pm 5$ $2.5\pm 0.2$ $8\pm 9.4$ $2.5\pm 0.4$ $0.3\pm 0.6$ $2.5\pm 0.2$

### v2

@@ I Introduction
A solution to increase collaboration is to give incentives in exchange for these data. This can be done in different ways, depending on whether the companies are willing to share data. If this is the case, different methods have been proposed for renewable power forecasting in the literature exploiting spatio-temporal relations. For solar power forecasting, a model incorporating exogenous data from different sources was developed [3], while a similar approach has been applied to both wind and solar power forecasting [1]. In wind power forecasting, sparse spatio-temporal models were introduced in [5] and later extended in an online learning context [11]. Furthermore, to avoid the loss of privacy companies could face, privacy-preserving solutions using distributed learning methods have been proposed [4], with further advancements addressing online settings [17, Gonçalves2021]. However, all these solutions assume that the agents are willing to collaborate to improve forecasts and that they act rationally and truthfully. This is not always the case in practice. An alternative approach is to view the problem in a more general framework, through prediction markets, in which companies choose to share their individual forecasts and get rewarded for their contribution to the resulting aggregate forecast.
@@
Nevertheless, existing solutions fail to consider some essential aspects of real-world applications such as: (i) real-time implementation, (ii) historical contributions of the participants, and (iii) the ability to accommodate intermittent participation, meaning that participants may join or exit the market at any time. The first two challenges can be seen as an online learning problem, whilst the third one as a missing data problem. For the latter, the literature has proposed robust variants of linear regression [2, 18] and online imputation strategies [9]. In addiction, a pay-off allocation mechanism must be designed to determine how rewards are distributed among sellers, reflecting the informational value they contribute to the final forecast. Also in this case, the allocation must reflect historical contributions rather than only instantaneous performance. In the literature, the most widely adopted approach is the Shapley-value-based allocation [15] (see, for example, [Gonçalves2021]) and online versions of Shapley value calculation have also been employed for dynamic pay-off distribution [13]. However, these approaches primarily focus on in-sample pay-offs and fail to account for out-of-sample performance, overlooking the actual contribution of sellers to genuine forecast improvements. Finally, any allocation rule must also satisfy several desirable economic properties to ensure fairness and stability.
@@ IV-B Performance with Varying Missingness
RQR $4.8\pm 3.9$ $1.7\pm 0.14$ $5\pm 6.5$ $1.7\pm 0.2$ $0.3\pm 4$ $1.7\pm 0.16$
Last Impute $5\pm 4.6$ $2.5\pm 0.2$ $35\pm 7$ $2.5\pm 0.3$ $18\pm 6$ $2.5\pm 0.2$
Mean Impute $7\pm 5$ $2.5\pm 0.2$ $8\pm 9.4$ $2.5\pm 0.4$ $0.3\pm 0.6$ $2.5\pm 0.2$